OpenStudy (anonymous):

Karen is eleven years less than four times John’s age. If Karen is 49, how old is John?

OpenStudy (goldphenoix):

49= 4j-11

OpenStudy (ivettef365):

you do the equation x = john age then you have 4x - 11 = 49 solve for x

OpenStudy (anonymous):

solve and ill give medal

OpenStudy (goldphenoix):

14.75 years old.

OpenStudy (anonymous):

??

OpenStudy (goldphenoix):

John is 14 years old.

OpenStudy (anonymous):

not an option

OpenStudy (goldphenoix):

He'll be 15 in 1/4 of a year.

OpenStudy (ivettef365):

no first you do 4x = 49 + 11 then you have 4x = 60 then divide 60/4

OpenStudy (ivettef365):

now John's age is 4x, so multiply 4 x 15

OpenStudy (goldphenoix):

My bad, you're right.

OpenStudy (ivettef365):

the question is what is John's age

OpenStudy (anonymous):

john is 60?